OpenStudy (anonymous):

At present time a man is four times as his son. 6 years ago he was 10 times as old. find their present age.

OpenStudy (anonymous):

Define some variables. Write two true equations.

OpenStudy (anonymous):

we can write this as a system of equations x = 4y x - 6 =10y x = dad's age y = son's age

OpenStudy (anonymous):

i messed up the second equation, it should be x - 6 = 10(y - 6)

OpenStudy (anonymous):

x=4y x-6=10(y-6)

OpenStudy (anonymous):

since x = 4y that means x - 6 = 10(y - 6) can also be written as 4y - 6 = 10(y - 6) solve for y 4y - 6 = 10y - 60 54 = 6y y = 9

OpenStudy (anonymous):

yeh x=4y x=10y-54 4y=10y-54............6y=54.......y=9......x=36

OpenStudy (anonymous):

if the son is 9 years old, according to our first equation x = 4y the dad should be 4(9) years old or 36 years old